Consider the boundary value problem(BVP) defined by laplaces equation.

uxx+uyy=0 on the square o<x,y<2

subjected to the boundary conditions:

u(0,y)=0 ,u(x,2)=0 ,u(2,y)=0 ,u(x,0)=100sin(πx/2)

Solve the BVP using method of separation of variable.


1
Expert's answer
2021-07-26T15:56:38-0400

The given equation is -


"\dfrac{\partial^{2}u}{\partial x^{2}}" "+\dfrac{\partial^{2}u}{\partial y^{2}}=0.........................................1)"


Let "u=XY.............................................2)"


where "X" is a function of "x" and "Y" is a function of "y" , only then ,


"\dfrac{\partial^{2}u}{\partial x^{2}}=\dfrac{\partial^{2}(XY)}{\partial x^{2}}=Y\dfrac{d^{2}X}{dx^{2}}" and


"\dfrac{\partial^{2}u}{\partial y^{2}}=\dfrac{\partial^{2}(XY)}{\partial y^{2}}=X\dfrac{d^{2}Y}{dy^{2}}"



"\therefore" From (1), "YX^{''}+XY^{''}=0"


"\implies" "\dfrac{X^{''}}{X}+\dfrac{Y^{''}}{Y}=0"


Now , we will make some cases -



Case I. When "\dfrac{X^{''}}{X}=-\dfrac{Y^{''}}{Y}" "=p^{2}(say)"


"i) \ \dfrac{X^{''}}{X}=p^{2}"


"X^{''}-P^{2}X=0"


Auxiliary equation is-


"m^{2}-p^{2}=0"


"m={\pm}p"


"\therefore" "CF=c_1e^{px}+c_2e^{-px}"


"PI=0"


"\therefore" "X=" "c_1e^{px}+c_2e^{-px}"


"ii)" "\ \dfrac{Y^{''}}{Y}=p^{2}\implies" "Y^{''}+P^{2}Y=0"


Auxiliary equation is "-"


"m^{2}+p^{2}=0\implies" "m={\pm}pi"


"\therefore" "CF=c_3cospy+c_4sinpy"


"PI=0"


"\therefore" "y=c_3cospy+c_4sinpy"


Now , we have given some boundary conditions making use of them now -


"X(0)=0"


"\implies" "c_{1}+c_{2}=0" "\implies c_2=-c_1"


"X(x)=0"


"\implies" "c_{1}e^{px}+c_2e^{-px}=0" "\implies" "c_1(e^{px}+e^{-px})=0"


"\implies" "c_1=0" (Since "e^{px}-e^{-px}" "\neq0(as\ p\neq0\neq l)"


"\therefore" "c_2=0"


"\therefore" "X=0\implies" "u=XY=0" which is impossible .


Hence Case 1 is rejected .


"Case \ 2)" when "\dfrac{X^{''}}{X}=-\dfrac{Y^{''}}{Y}=0"


"i)" "\dfrac{X^{''}}{X}=0"


"\implies" "X^{''}=0\implies" "X=c_5x+c_6"


"ii)" "-\dfrac{Y^{''}}{Y}=0"


"\implies" "Y^{''}=0\implies" "Y=c_7y+c_8"


Now , "X(0)=0\implies" "c_6" "=" 0


"X(x)=0"


"\implies" "c_5+c_6=0\implies""c_5l=0"


"\implies" "c_5=0"


"\therefore X=0"

"\therefore u=XY=0" which is impossible .


Hence we reject case 2 also .


"Case \ 3)" when


"\dfrac{X^{''}}{X}=-\dfrac{Y^{''}}{Y}=-p^{2}(say)"



"(i)" "\dfrac{X^{''}}{X}=-p^{2}" "\implies \dfrac {d^{2}X}{dx^{2}}+p^{2}=0"


"\implies" "X^{''}+p^{2}X=0"


Auxiliary equation is given by -


"=m^{2}+p^{2}=0\implies" "m={\pm}pi"


"CF=c_9cospx+c_{10}sinpx"


"PI=0"


"X=c_9cospx+c_{10}sinpx"


"(ii)" "-\dfrac{Y^{''}}{Y}=-p^{2}" "\implies \dfrac{d^{2}y}{dy^{2}}-p^{2}y=0"


Auxiliary equation is -


"m^{2}-p^{2}=0"


"m={\pm}p"


"\therefore CF=c_{11}e^{py}+c_{12}e^{-py}"


"PI=0"


hence , "Y=c_{11}e^{py}+c_{12}e^{-py}"


Now , "X(0)=0 \implies c_{9}=0"


"\therefore X=c_{10}sinpx"


"X(l)=0"


"=c_{10}sinpl=0"


"\therefore sinpl=0=sinn{\pi}, n\in I"


"\therefore" "p=\dfrac{n{\pi}}l"


"\therefore" "X=c_{10}sin\dfrac{n{\pi}x}{l}" .............................................................3)


Again "Y(0)=0"


"\therefore c_{11}+c_{12}=0 \therefore c_{11}=-c_{12}"



"Y=c_{11}(e^{py}-e^{-py})=c_{11}(e^\dfrac{n{\pi}}{l}-e^\dfrac{-n{\pi}}{l})"



"\therefore u=XY=c_{10}c_{11}sin\dfrac{n{\pi}x}{l}[(e^\dfrac{n{\pi}}{l}-e^\dfrac{-n{\pi}}{l})]"



or "u(x,y)=b_nsin\dfrac{n{\pi}x}{l}[(e^\dfrac{n{\pi}}{l}-e^\dfrac{-n{\pi}}{l})]"



Now "u(x,0)=100sin\dfrac{{\pi}x}{2}=b_nsin\dfrac{n{\pi}x}{l}[(e^\dfrac{n{\pi}}{l}-e^\dfrac{-n{\pi}}{l})]"


"=b_{n}=\dfrac{100sin\dfrac{{\pi}x}{2}}{sin\dfrac{n{\pi}x}{l}[(e^\dfrac{n{\pi}}{l}-e^\dfrac{-n{\pi}}{l})]}"



"u(x,y)=" "[(e^\dfrac{n{\pi}}{l}-e^\dfrac{-n{\pi}}{l})]\dfrac{100sin\dfrac{{\pi}x}{2}}{sin\dfrac{n{\pi}x}{l}[(e^\dfrac{n{\pi}}{l}-e^\dfrac{-n{\pi}}{l})]}" "sin\dfrac{n{\pi}x}{l}"

"u(x,y)=100sin\dfrac{{\pi}x}{2}"


which is given solution .